OpenStudy (anonymous):

A triangle with side lengths of 5, 12, and 13 is a right triangle. I think it is. Is that correct guys?

OpenStudy (133):

A right angle is 90 degrees.

OpenStudy (anonymous):

\[ 13^2=169= 12^2 + 5^2=169 \] So it is a right triangle